# A Survey of Protocol Fuzzing

### v2

@@ Acknowledgements.
- Jero et al. (2018) Samuel Jero, Md. Endadul Hoque, David R. Choffnes, Alan Mislove, and Cristina Nita-Rotaru. 2018. Automated Attack Discovery in TCP Congestion Control Using a Model-guided Approach. In 25th Annual Network and Distributed System Security Symposium, NDSS 2018, San Diego, California, USA, February 18-21, 2018. The Internet Society. http://wp.internetsociety.org/ndss/wp-content/uploads/sites/25/2018/02/ndss2018_02A-1_Jero_paper.pdf
@@
- Jero et al. (2015) Samuel Jero, Hyojeong Lee, and Cristina Nita-Rotaru. 2015. Leveraging State Information for Automated Attack Discovery in Transport Protocol Implementations. In 2015 45th Annual IEEE/IFIP International Conference on Dependable Systems and Networks. 1–12. https://doi.org/10.1109/DSN.2015.22
@@
- Jero et al. (2019) Samuel Jero, Maria Leonor Pacheco, Dan Goldwasser, and Cristina Nita-Rotaru. 2019. Leveraging Textual Specifications for Grammar-Based Fuzzing of Network Protocols. Proceedings of the AAAI Conference on Artificial Intelligence 33, 01 (Jul. 2019), 9478–9483. https://doi.org/10.1609/aaai.v33i01.33019478
@@
- Ji and Xu (2023) Ru Ji and Meng Xu. 2023. Finding Specification Blind Spots via Fuzz Testing. In 2023 IEEE Symposium on Security and Privacy (SP). IEEE Computer Society, 2708–2725.
@@
- Jung et al. (2001) Jaeyeon Jung, Emil Sit, Hari Balakrishnan, and Robert Morris. 2001. DNS performance and the effectiveness of caching. In Proceedings of the 1st ACM SIGCOMM Workshop on Internet Measurement. 153–167.
@@
- Jurjens (2006) Jan Jurjens. 2006. Security Analysis of Crypto-based Java Programs using Automated Theorem Provers. In 21st IEEE/ACM International Conference on Automated Software Engineering (ASE’06). 167–176. https://doi.org/10.1109/ASE.2006.60
@@
- Katz and Shin (2006) Jonathan Katz and Ji Sun Shin. 2006. Parallel and concurrent security of the HB and HB+ protocols. In Advances in Cryptology - EUROCRYPT 2006. Springer, 73–87.
@@
- Kim et al. (2017) Seulbae Kim, Seunghoon Woo, Heejo Lee, and Hakjoo Oh. 2017. Poster: Iotcube: an automated analysis platform for finding security vulnerabilities. In Proceedings of the 38th IEEE Symposium on Poster presented at Security and Privacy.
@@
- Krupp et al. (2022) Johannes Krupp, Ilya Grishchenko, and Christian Rossow. 2022. AmpFuzz: Fuzzing for Amplification DDoS Vulnerabilities. In 31st USENIX Security Symposium (USENIX Security 22). USENIX Association, Boston, MA, 1043–1060. https://www.usenix.org/conference/usenixsecurity22/presentation/krupp
@@
- Lee et al. (2018) Seungsoo Lee, Jinwoo Kim, Seungwon Woo, and Seungwon Shin. 2018. The Finest Penetration Testing Framework for Software-Defined Networks. https://www.blackhat.com/us-18/briefings/schedule/#the-finest-penetration-testing-framework-for-software-defined-networks--10101
@@
- Li et al. (2022b) Ao Li, Rohan Padhye, and Vyas Sekar. 2022b. SPIDER: A Practical Fuzzing Framework to Uncover Stateful Performance Issues in SDN Controllers. https://doi.org/10.48550/ARXIV.2209.04026
@@
- Li et al. (2022a) Junqiang Li, Senyi Li, Gang Sun, Ting Chen, and Hongfang Yu. 2022a. SNPSFuzzer: A Fast Greybox Fuzzer for Stateful Network Protocols using Snapshots. CoRR abs/2202.03643 (2022). arXiv:2202.03643 https://arxiv.org/abs/2202.03643
- Li et al. (2020) Yuekang Li, Hongxu Chen, Cen Zhang, Siyang Xiong, Chaoyi Liu, and Yi Wang. 2020. Ori: A greybox fuzzer for SOME/IP protocols in automotive Ethernet. In 2020 27th Asia-Pacific Software Engineering Conference (APSEC). IEEE, 495–499.
- Li et al. (2021) Yuekang Li, Guozhu Meng, Jun Xu, Cen Zhang, Hongxu Chen, Xiaofei Xie, Haijun Wang, and Yang Liu. 2021. Vall-nut: Principled Anti-Grey box-Fuzzing. In 2021 IEEE 32nd International Symposium on Software Reliability Engineering (ISSRE). IEEE, 288–299.
- Li et al. (2019) Yuekang Li, Yinxing Xue, Hongxu Chen, Xiuheng Wu, Cen Zhang, Xiaofei Xie, Haijun Wang, and Yang Liu. 2019. Cerebro: context-aware adaptive fuzzing for effective vulnerability detection. In Proceedings of the 2019 27th ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering. 533–544.
@@
- Liang et al. (2018) Hongliang Liang, Xiaoxiao Pei, Xiaodong Jia, Wuwei Shen, and Jian Zhang. 2018. Fuzzing: State of the Art. IEEE Transactions on Reliability 67, 3 (2018), 1199–1218. https://doi.org/10.1109/TR.2018.2834476
@@
- Liu et al. (2020) Dongge Liu, Gidon Ernst, Toby Murray, and Benjamin I. P. Rubinstein. 2020. Legion: Best-First Concolic Testing. In Proceedings of the 35th IEEE/ACM International Conference on Automated Software Engineering (Virtual Event, Australia) (ASE ’20). Association for Computing Machinery, New York, NY, USA, 54–65. https://doi.org/10.1145/3324884.3416629
@@
- Liu et al. (2021) Dongge Liu, Van-Thuan Pham, Gidon Ernst, Toby Murray, and Benjamin I. P. Rubinstein. 2021. State Selection Algorithms and Their Impact on The Performance of Stateful Network Protocol Fuzzing. CoRR abs/2112.15498 (2021). arXiv:2112.15498 https://arxiv.org/abs/2112.15498
@@
- Liu et al. (2022) Puzhuo Liu, Yaowen Zheng, Zhanwei Song, Dongliang Fang, Shichao Lv, and Limin Sun. 2022. Fuzzing proprietary protocols of programmable controllers to find vulnerabilities that affect physical control. Journal of Systems Architecture 127 (2022), 102483.
@@
- Luo et al. (2023) Zhengxiong Luo, Junze Yu, Feilong Zuo, Jianzhong Liu, Yu Jiang, Ting Chen, Abhik Roychoudhury, and Jiaguang Sun. 2023. Bleem: Packet Sequence Oriented Fuzzing for Protocol Implementations. In 32nd USENIX Security Symposium (USENIX Security 23). USENIX Association, Anaheim, CA, 4481–4498. https://www.usenix.org/conference/usenixsecurity23/presentation/luo-zhengxiong
@@
- Luo et al. (2019) Zhengxiong Luo, Feilong Zuo, Yu Jiang, Jian Gao, Xun Jiao, and Jiaguang Sun. 2019. Polar: Function Code Aware Fuzz Testing of ICS Protocol. ACM Trans. Embed. Comput. Syst. 18, 5s (2019), 93:1–93:22. https://doi.org/10.1145/3358227
@@
- Luo et al. (2020) Zhengxiong Luo, Feilong Zuo, Yuheng Shen, Xun Jiao, Wanli Chang, and Yu Jiang. 2020. ICS Protocol Fuzzing: Coverage Guided Packet Crack and Generation. In 57th ACM/IEEE Design Automation Conference, DAC 2020, San Francisco, CA, USA, July 20-24, 2020. IEEE, 1–6. https://doi.org/10.1109/DAC18072.2020.9218603
@@
- Ma et al. (2023) Fuchen Ma, Yuanliang Chen, Meng Ren, Yuanhang Zhou, Yu Jiang, Ting Chen, Huizhong Li, and Jiaguang Sun. 2023. LOKI: State-Aware Fuzzing Framework for the Implementation of Blockchain Consensus Protocols. In Proceedings 2023 Network and Distributed System Security Symposium.
@@
- Maier et al. (2020) Dominik Maier, Lukas Seidel, and Shinjo Park. 2020. BaseSAFE: Baseband SAnitized Fuzzing through Emulation. CoRR abs/2005.07797 (2020). arXiv:2005.07797 https://arxiv.org/abs/2005.07797
@@
- Manes et al. (2021) V. M. Manes, H. Han, C. Han, S. Cha, M. Egele, E. J. Schwartz, and M. Woo. 2021. The Art, Science, and Engineering of Fuzzing: A Survey. IEEE Transactions on Software Engineering 47, 11 (nov 2021), 2312–2331. https://doi.org/10.1109/TSE.2019.2946563
@@
- Marsden (1986) B.W. Marsden. 1986. Communication Network Protocols. Chartwell-Bratt. https://books.google.com.hk/books?id=WEeOzgEACAAJ
- Mathis et al. (2019) Björn Mathis, Rahul Gopinath, Michaël Mera, Alexander Kampmann, Matthias Höschele, and Andreas Zeller. 2019. Parser-Directed Fuzzing. In Proceedings of the 40th ACM SIGPLAN Conference on Programming Language Design and Implementation (Phoenix, AZ, USA) (PLDI 2019). Association for Computing Machinery, New York, NY, USA, 548–560. https://doi.org/10.1145/3314221.3314651
- McMahon Stone et al. (2018) Chris McMahon Stone, Tom Chothia, and Joeri de Ruiter. 2018. Extending Automated Protocol State Learning for the 802.11 4-Way Handshake. In Computer Security, Javier Lopez, Jianying Zhou, and Miguel Soriano (Eds.). Springer International Publishing, Cham, 325–345.
- Meng et al. (2021) Ruijie Meng, Zhen Dong, Jialin Li, Ivan Beschastnikh, and Abhik Roychoudhury. 2021. Finding Counterexamples of Temporal Logic properties in Software Implementations via Greybox Fuzzing. CoRR abs/2109.02312 (2021). arXiv:2109.02312 https://arxiv.org/abs/2109.02312
- Meng et al. (2024) Ruijie Meng, Martin Mirchev, Marcel Böhme, and Abhik Roychoudhury. 2024. Large Language Model guided Protocol Fuzzing. In NDSS. 1–17.
- Meng et al. (2023) Ruijie Meng, George Pîrlea, Abhik Roychoudhury, and Ilya Sergey. 2023. Greybox Fuzzing of Distributed Systems. In Proceedings of the 2023 ACM SIGSAC Conference on Computer and Communications Security (¡conf-loc¿, ¡city¿Copenhagen¡/city¿, ¡country¿Denmark¡/country¿, ¡/conf-loc¿) (CCS ’23). Association for Computing Machinery, New York, NY, USA, 1615–1629. https://doi.org/10.1145/3576915.3623097
@@
- Musuvathi and Engler (2004) Madanlal Musuvathi and Dawson R. Engler. 2004. Model Checking Large Network Protocol Implementations. In Proceedings of the 1st Conference on Symposium on Networked Systems Design and Implementation - Volume 1 (San Francisco, California) (NSDI’04). USENIX Association, USA, 12.
@@
- Ohkawa et al. (2019) Takeshi Ohkawa, Yuhei Sugata, Harumi Watanabe, Nobuhiko Ogura, Kanemitsu Ootsu, and Takashi Yokota. 2019. High level synthesis of ROS protocol interpretation and communication circuit for FPGA. In 2019 IEEE/ACM 2nd International Workshop on Robotics Software Engineering (RoSE). IEEE, 33–36.
@@
- Ozavci (2013) Fatih Ozavci. 2013. VoIP Wars : Return of the SIP. https://media.defcon.org/DEF%20CON%2021/DEF%20CON%2021%20presentations/DEF%20CON%2021%20-%20Ozavci-VoIP-Wars-Return-of-the-SIP.pdf
- Pacheco et al. (2022) Maria Leonor Pacheco, Max von Hippel, Ben Weintraub, Dan Goldwasser, and Cristina Nita-Rotaru. 2022. Automated Attack Synthesis by Extracting Finite State Machines from Protocol Specification Documents. CoRR abs/2202.09470 (2022). arXiv:2202.09470 https://arxiv.org/abs/2202.09470
@@
- Park et al. (2022) H. Park, C. Nkuba, S. Woo, and H. Lee. 2022. L2Fuzz: Discovering Bluetooth L2CAP Vulnerabilities Using Stateful Fuzz Testing. In 2022 52nd Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN). IEEE Computer Society, Los Alamitos, CA, USA, 343–354. https://doi.org/10.1109/DSN53405.2022.00043
- Pearson et al. (2022) Bryan Pearson, Yue Zhang, Cliff Zou, and Xinwen Fu. 2022. FUME: Fuzzing Message Queuing Telemetry Transport Brokers. In IEEE INFOCOM 2022 - IEEE Conference on Computer Communications. 1699–1708. https://doi.org/10.1109/INFOCOM48880.2022.9796755
- Peterson et al. (2020) Anthony Peterson, Samuel Jero, Endadul Hoque, David Choffnes, and Cristina Nita-Rotaru. 2020. aBBRate: Automating BBR Attack Exploration Using a Model-Based Approach. In 23rd International Symposium on Research in Attacks, Intrusions and Defenses (RAID 2020). USENIX Association, San Sebastian, 225–240. https://www.usenix.org/conference/raid2020/presentation/peterson
- Pham et al. (2020) Van-Thuan Pham, Marcel Böhme, and Abhik Roychoudhury. 2020. AFLNET: A Greybox Fuzzer for Network Protocols. In 13th IEEE International Conference on Software Testing, Validation and Verification, ICST 2020, Porto, Portugal, October 24-28, 2020. IEEE, 460–465. https://doi.org/10.1109/ICST46399.2020.00062
@@
- Poncelet et al. (2022) Clément Poncelet, Konstantinos Sagonas, and Nicolas Tsiftes. 2022. So Many Fuzzers, So Little Time: Experience from Evaluating Fuzzers on the Contiki-NG Network (Hay) Stack. In Proceedings of the 37th IEEE/ACM International Conference on Automated Software Engineering. 1–12.
@@
- Qu et al. (2021) Lewei Qu, Dongxiang Ke, Ye Zhang, and Ying Wang. 2021. BadMesher: New Attack Surfaces of Wi-Fi Mesh Network. https://www.blackhat.com/eu-21/briefings/schedule/#badmesher-new-attack-surfaces-of-wi-fi-mesh-network-24181
@@
- Rangnau et al. (2020) Thorsten Rangnau, Remco v. Buijtenen, Frank Fransen, and Fatih Turkmen. 2020. Continuous Security Testing: A Case Study on Integrating Dynamic Security Testing Tools in CI/CD Pipelines. In 2020 IEEE 24th International Enterprise Distributed Object Computing Conference (EDOC). 145–154. https://doi.org/10.1109/EDOC49727.2020.00026
- Reen and Rossow (2020) Gaganjeet Singh Reen and Christian Rossow. 2020. DPIFuzz: A Differential Fuzzing Framework to Detect DPI Elusion Strategies for QUIC. In Annual Computer Security Applications Conference (Austin, USA) (ACSAC ’20). Association for Computing Machinery, New York, NY, USA, 332–344. https://doi.org/10.1145/3427228.3427662
- Ren et al. (2021) Mengfei Ren, Xiaolei Ren, Huadong Feng, Jiang Ming, and Yu Lei. 2021. Z-Fuzzer: device-agnostic fuzzing of Zigbee protocol implementation. In WiSec ’21: 14th ACM Conference on Security and Privacy in Wireless and Mobile Networks, Abu Dhabi, United Arab Emirates, 28 June - 2 July, 2021, Christina Pöpper, Mathy Vanhoef, Lejla Batina, and René Mayrhofer (Eds.). ACM, 347–358. https://doi.org/10.1145/3448300.3468296

### v3

@@ Acknowledgements.
- Jero et al. (2018) Samuel Jero, Md. Endadul Hoque, David R. Choffnes, Alan Mislove, and Cristina Nita-Rotaru. 2018. Automated Attack Discovery in TCP Congestion Control Using a Model-guided Approach. In 25th Annual Network and Distributed System Security Symposium (NDSS).
@@
- Jero et al. (2015) Samuel Jero, Hyojeong Lee, and Cristina Nita-Rotaru. 2015. Leveraging State Information for Automated Attack Discovery in Transport Protocol Implementations. In 2015 45th Annual IEEE/IFIP International Conference on Dependable Systems and Networks. 1–12.
@@
- Jero et al. (2019) Samuel Jero, Maria Leonor Pacheco, Dan Goldwasser, and Cristina Nita-Rotaru. 2019. Leveraging Textual Specifications for Grammar-Based Fuzzing of Network Protocols. AAAI Conference on Artificial Intelligence 33, 01 (2019), 9478–9483. https://ojs.aaai.org/index.php/AAAI/article/view/5002
@@
- Ji and Xu (2023) Ru Ji and Meng Xu. 2023. Finding Specification Blind Spots via Fuzz Testing. In 2023 IEEE Symposium on Security and Privacy (SP). 2708–2725.
@@
- Jung et al. (2001) Jaeyeon Jung, Emil Sit, Hari Balakrishnan, and Robert Morris. 2001. DNS performance and the effectiveness of caching. In 1st ACM SIGCOMM Workshop on Internet Measurement. 153–167.
@@
- Jurjens (2006) Jan Jurjens. 2006. Security Analysis of Crypto-based Java Programs using Automated Theorem Provers. In 21st IEEE/ACM International Conference on Automated Software Engineering (ASE’06). 167–176.
- Kakarla and Beckett (2023) Siva Kesava Reddy Kakarla and Ryan Beckett. 2023. Oracle-based Protocol Testing with Eywa. arXiv:2312.06875 (2023).
@@
- Katz and Shin (2006) Jonathan Katz and Ji Sun Shin. 2006. Parallel and concurrent security of the HB and HB+ protocols. In 25th International Cryptology Conference (EUROCRYPT). 73–87.
@@
- Kim et al. (2017) Seulbae Kim, Seunghoon Woo, Heejo Lee, and Hakjoo Oh. 2017. Poster: Iotcube: an automated analysis platform for finding security vulnerabilities. In 38th IEEE Symposium on Poster presented at Security and Privacy.
@@
- Krupp et al. (2022) Johannes Krupp, Ilya Grishchenko, and Christian Rossow. 2022. AmpFuzz: Fuzzing for Amplification DDoS Vulnerabilities. In 31st USENIX Security Symposium (USENIX Security 22). 1043–1060.
@@
- Lee et al. (2018) Seungsoo Lee, Jinwoo Kim, Seungwon Woo, and Seungwon Shin. 2018. The Finest Penetration Testing Framework for Software-Defined Networks. In Blackhat US 2018.
@@
- Li et al. (2022b) Ao Li, Rohan Padhye, and Vyas Sekar. 2022b. SPIDER: A Practical Fuzzing Framework to Uncover Stateful Performance Issues in SDN Controllers. https://arxiv.org/abs/2209.04026
@@
- Li et al. (2022a) Junqiang Li, Senyi Li, Gang Sun, Ting Chen, and Hongfang Yu. 2022a. SNPSFuzzer: A Fast Greybox Fuzzer for Stateful Network Protocols using Snapshots. (2022). arXiv:2202.03643
@@
- Liang et al. (2018) Hongliang Liang, Xiaoxiao Pei, Xiaodong Jia, Wuwei Shen, and Jian Zhang. 2018. Fuzzing: State of the Art. IEEE Transactions on Reliability 67, 3 (2018), 1199–1218.
@@
- Liu et al. (2020) Dongge Liu, Gidon Ernst, Toby Murray, and Benjamin I. P. Rubinstein. 2020. Legion: Best-First Concolic Testing. In 35th IEEE/ACM International Conference on Automated Software Engineering (ASE). 54–65.
@@
- Liu et al. (2022a) D. Liu, V. Pham, G. Ernst, T. Murray, and B. P. Rubinstein. 2022a. State Selection Algorithms and Their Impact on The Performance of Stateful Network Protocol Fuzzing. In 2022 IEEE International Conference on Software Analysis, Evolution and Reengineering (SANER). 720–730.
@@
- Liu et al. (2022b) Puzhuo Liu, Yaowen Zheng, Zhanwei Song, Dongliang Fang, Shichao Lv, and Limin Sun. 2022b. Fuzzing proprietary protocols of programmable controllers to find vulnerabilities that affect physical control. Journal of Systems Architecture 127 (2022), 102483.
- Liu et al. (2021) Qiang Liu, Cen Zhang, Lin Ma, Muhui Jiang, Yajin Zhou, Lei Wu, Wenbo Shen, Xiapu Luo, Yang Liu, and Kui Ren. 2021. FirmGuide: Boosting the Capability of Rehosting Embedded Linux Kernels through Model-Guided Kernel Execution. In 36th IEEE/ACM International Conference on Automated Software Engineering (ASE). 792–804.
- Luo et al. (2024a) Zhengxiong Luo, Kai Liang, Yanyang Zhao, Feifan Wu, Junze Yu, Heyuan Shi, and Yu Jiang. 2024a. DYNPRE: Protocol Reverse Engineering via Dynamic Inference. In 31th Annual Network and Distributed System Security Symposium (NDSS). 1–18.
- Luo et al. (2024b) Zhengxiong Luo, Junze Yu, Qingpeng Du, Yanyang Zhao, Feifan Wu, and Heyuan Shi. 2024b. Parallel Fuzzing of IoT Messaging Protocols through Collaborative Packet Generation. In ACM SIGBED International Conference on Embedded Software (EMSOFT). 1–12.
@@
- Luo et al. (2023) Zhengxiong Luo, Junze Yu, Feilong Zuo, Jianzhong Liu, Yu Jiang, Ting Chen, Abhik Roychoudhury, and Jiaguang Sun. 2023. Bleem: Packet Sequence Oriented Fuzzing for Protocol Implementations. In 32nd USENIX Security Symposium (USENIX Security 23). 4481–4498.
@@
- Luo et al. (2019) Zhengxiong Luo, Feilong Zuo, Yu Jiang, Jian Gao, Xun Jiao, and Jiaguang Sun. 2019. Polar: Function Code Aware Fuzz Testing of ICS Protocol. ACM Trans. Embed. Comput. Syst. 18, 5s (2019), 93:1–93:22.
@@
- Luo et al. (2020) Zhengxiong Luo, Feilong Zuo, Yuheng Shen, Xun Jiao, Wanli Chang, and Yu Jiang. 2020. ICS Protocol Fuzzing: Coverage Guided Packet Crack and Generation. In 57th ACM/IEEE Design Automation Conference (DAC). 1–6.
@@
- Ma et al. (2023a) Fuchen Ma, Yuanliang Chen, Meng Ren, Yuanhang Zhou, Yu Jiang, Ting Chen, Huizhong Li, and Jiaguang Sun. 2023a. LOKI: State-Aware Fuzzing Framework for the Implementation of Blockchain Consensus Protocols. In Proceedings 2023 Network and Distributed System Security Symposium.
- Ma et al. (2024) Xiaoyue Ma, Lannan Luo, and Qiang Zeng. 2024. From One Thousand Pages of Specification to Unveiling Hidden Bugs: Large Language Model Assisted Fuzzing of Matter IoT Devices. In 33rd USENIX Security Symposium (USENIX Security 24). 4783–4800.
- Ma et al. (2023b) Xiaoyue Ma, Qiang Zeng, Haotian Chi, and Lannan Luo. 2023b. No more companion apps hacking but one dongle: Hub-based blackbox fuzzing of iot firmware. In 21st Annual International Conference on Mobile Systems, Applications and Services (Mobisys). 205–218.
@@
- Maier et al. (2020) Dominik Maier, Lukas Seidel, and Shinjo Park. 2020. BaseSAFE: Baseband SAnitized Fuzzing through Emulation. (2020). arXiv:2005.07797
@@
- Manes et al. (2021) V. M. Manes, H. Han, C. Han, S. Cha, M. Egele, E. J. Schwartz, and M. Woo. 2021. The Art, Science, and Engineering of Fuzzing: A Survey. IEEE Transactions on Software Engineering 47, 11 (2021), 2312–2331.
@@
- Marsden (1986) B.W. Marsden. 1986. Communication Network Protocols.
- Mathis et al. (2019) Björn Mathis, Rahul Gopinath, Michaël Mera, Alexander Kampmann, Matthias Höschele, and Andreas Zeller. 2019. Parser-Directed Fuzzing. In 40th ACM SIGPLAN Conference on Programming Language Design and Implementation (PLDI). 548–560.
- McMahon Stone et al. (2018) Chris McMahon Stone, Tom Chothia, and Joeri de Ruiter. 2018. Extending Automated Protocol State Learning for the 802.11 4-Way Handshake. In Computer Security. 325–345.
- McMahon Stone et al. (2022) Chris McMahon Stone, Sam L. Thomas, Mathy Vanhoef, James Henderson, Nicolas Bailluet, and Tom Chothia. 2022. The Closer You Look, The More You Learn: A Grey-box Approach to Protocol State Machine Learning. In 2022 ACM SIGSAC Conference on Computer and Communications Security (CCS). 2265–2278.
- Meng et al. (2021) Ruijie Meng, Zhen Dong, Jialin Li, Ivan Beschastnikh, and Abhik Roychoudhury. 2021. Finding Counterexamples of Temporal Logic properties in Software Implementations via Greybox Fuzzing. CoRR abs/2109.02312 (2021). arXiv:2109.02312
- Meng et al. (2024) Ruijie Meng, Martin Mirchev, Marcel Böhme, and Abhik Roychoudhury. 2024. Large Language Model guided Protocol Fuzzing. In 31st Annual Network and Distributed System Security Symposium (NDSS). 1–17.
- Meng et al. (2023) Ruijie Meng, George Pîrlea, Abhik Roychoudhury, and Ilya Sergey. 2023. Greybox Fuzzing of Distributed Systems. In 2023 ACM SIGSAC Conference on Computer and Communications Security (CCS). 1615–1629.
@@
- Musuvathi and Engler (2004) Madanlal Musuvathi and Dawson R. Engler. 2004. Model Checking Large Network Protocol Implementations. In First Symposium on Networked Systems Design and Implementation (NSDI 04). 12.
@@
- Nikbakht et al. (2024) Rasoul Nikbakht, Mohamed Benzaghta, and Giovanni Geraci. 2024. TSpec-LLM: An Open-source Dataset for LLM Understanding of 3GPP Specifications. arXiv:2406.01768 (2024).
- NOCHVAY (2019) NOCHVAY. 2019. Security research: CODESYS Runtime, a PLC control framework. https://ics-cert.kaspersky.com/publications/reports/2019/09/18/security-research-codesys-runtime-a-plc-control-framework-part-1/
@@
- Ohkawa et al. (2019) Takeshi Ohkawa, Yuhei Sugata, Harumi Watanabe, Nobuhiko Ogura, Kanemitsu Ootsu, and Takashi Yokota. 2019. High level synthesis of ROS protocol interpretation and communication circuit for FPGA. In 2nd International Workshop on Robotics Software Engineering (RoSE). 33–36.
@@
- openvpn (2014) openvpn. 2014. OpenVPN: OpenVPN source code documentation. https://build.openvpn.net/doxygen/
- Ozavci (2013) Fatih Ozavci. 2013. VoIP Wars : Return of the SIP. In DEFCON 21.
- Pacheco et al. (2022) Maria Leonor Pacheco, Max von Hippel, Ben Weintraub, Dan Goldwasser, and Cristina Nita-Rotaru. 2022. Automated Attack Synthesis by Extracting Finite State Machines from Protocol Specification Documents. (2022). arXiv:2202.09470
@@
- Park et al. (2022) H. Park, C. Nkuba, S. Woo, and H. Lee. 2022. L2Fuzz: Discovering Bluetooth L2CAP Vulnerabilities Using Stateful Fuzz Testing. In 2022 52nd Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN). 343–354.
- Pearson et al. (2022) Bryan Pearson, Yue Zhang, Cliff Zou, and Xinwen Fu. 2022. FUME: Fuzzing Message Queuing Telemetry Transport Brokers. In 2022 IEEE Conference on Computer Communications (INFOCOM). 1699–1708.
- Peterson et al. (2020) Anthony Peterson, Samuel Jero, Endadul Hoque, David Choffnes, and Cristina Nita-Rotaru. 2020. aBBRate: Automating BBR Attack Exploration Using a Model-Based Approach. In 23rd International Symposium on Research in Attacks, Intrusions and Defenses (RAID). 225–240.
- Pham et al. (2020) Van-Thuan Pham, Marcel Böhme, and Abhik Roychoudhury. 2020. AFLNET: A Greybox Fuzzer for Network Protocols. In 13th IEEE International Conference on Software Testing, Validation and Verification (ICST). 460–465.
@@
- Poncelet et al. (2022) Clément Poncelet, Konstantinos Sagonas, and Nicolas Tsiftes. 2022. So Many Fuzzers, So Little Time: Experience from Evaluating Fuzzers on the Contiki-NG Network (Hay) Stack. In 37th IEEE/ACM International Conference on Automated Software Engineering. 1–12.
@@
- Qu et al. (2021) Lewei Qu, Dongxiang Ke, Ye Zhang, and Ying Wang. 2021. BadMesher: New Attack Surfaces of Wi-Fi Mesh Network. In Blackhat EU 2021.
@@
- Rangnau et al. (2020) Thorsten Rangnau, Remco v. Buijtenen, Frank Fransen, and Fatih Turkmen. 2020. Continuous Security Testing: A Case Study on Integrating Dynamic Security Testing Tools in CI/CD Pipelines. In 24th International Enterprise Distributed Object Computing Conference (EDOC). 145–154.
- Reen and Rossow (2020) Gaganjeet Singh Reen and Christian Rossow. 2020. DPIFuzz: A Differential Fuzzing Framework to Detect DPI Elusion Strategies for QUIC. In Annual Computer Security Applications Conference (ACSAC). 332–344.
- Ren et al. (2021) Mengfei Ren, Xiaolei Ren, Huadong Feng, Jiang Ming, and Yu Lei. 2021. Z-Fuzzer: device-agnostic fuzzing of Zigbee protocol implementation. In 14th ACM Conference on Security and Privacy in Wireless and Mobile Networks (WiSec). 347–358.
- Ren et al. (2023) Mengfei Ren, Haotian Zhang, Xiaolei Ren, Jiang Ming, and Yu Lei. 2023. Intelligent Zigbee Protocol Fuzzing via Constraint-Field Dependency Inference. In European Symposium on Research in Computer Security (ESORICS). 467–486.